# Electron polarization induced by electroweak interaction in high-energy scattering off light nuclei

### v1

@@ Abstract
Electron polarization plays a significant role in studies on nuclear scattering. Nevertheless, the development of a comprehensive approach to such a problem remains challenging, particularly at the relativistic electron-energy scale. Herein, we present a theoretical approach to investigate the impact of electron polarization in scattering off unoriented light nuclei, based on the multipole expansion for the scattering cross section at high energies. Numerical calculations for stable 6,7Li and unstable 7Be nuclei show that the longitudinal polarization and weak interaction are not explicitly correlated when electrons scatter at 0 for all energy scales. In contrast, their correlation is strongly exhibited at other scattering angles when energy exceeds 10 GeV. A comparison between stable and unstable nuclei reveals that the stability of nuclei significantly affects the electron polarization contribution to the scattering cross section. This study opens new approaches to the nuclear structure problems, in particular the EMC effect, via the deep inelastic electron-nucleus scattering using the unified electroweak theory.
@@
$\sigma_{\text{total}}^{\text{pol.}}(\xi,\xi^{\prime})=\sigma_{\text{total}}^{0}+\sigma_{\text{total}}^{\xi\xi^{\prime}},\vspace{-20pt}$ (1)
where $\xi$ and $\xi^{\prime}$ are the electron polarizations before and after scattering, whose values are equal to +1 or $-1$. The first term in Eq. (1), which does not involve electron polarization, is a sum of
@@
Obviously, the polarized cross sections [Eqs. (3a)–(3c)] depend on the polarization degrees $\xi$ and $\xi^{\prime}$. For simplicity, we assign $\xi\,(\xi^{\prime})=+1$ for the right-handed polarization and $\xi\,(\xi^{\prime})=-1$ for the left-handed one. We then consider all possible variations of electron polarization as following cases. Case 1: both incoming and outgoing electron beams are unpolarized. This case has been studied in detail in our recent study [14]. Case 2: electron polarizations before and after scattering are preserved, i.e., $\xi=\xi^{\prime}$. Case 3: an initially polarized beam becomes unpolarized, i.e., $\xi=+1/-1$ and $\xi^{\prime}=0$ (or $\xi^{\prime}$ is taken over +1 and $-1$). Case 4: an electron beam reverses its initial polarization after scattering, i.e., $\xi=-\xi^{\prime}$.
@@
Fig. 1 shows the calculated $\sigma_{\text{total}}^{\text{pol.}}(1,0)/\sigma_{\text{total}}^{\text{unpol.}}$ ratio for 6Li, whereas Fig. 2 plots the computed $\sigma_{\text{total}}^{\text{pol.}}(1,1)/\sigma_{\text{total}}^{\text{unpol.}}$ one for 7Li. The $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ ratio for 7Be is also displayed in Fig. 3. All the crosssection ratios for three considered nuclei are listed in Tab. I. It can be seen that at energies not exceeding $10$ GeV, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ is just half of $\sigma_{\text{total}}^{\text{unpol.}}$, while $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is almost equal to the latter. A slight difference is observed when comparing the cross-section ratios for three nuclei (Tab. I). This can be attributed to the negligible contribution of weak interaction to the polarized terms at such an energy range [13]. When the electron energy goes above 10 MeV, the weak interaction gradually emerges and correlates with electron polarization to make polarized cross sections considerably increase, as seen in Figs. 1–3.
@@
In conclusion, we have performed a systematic study on the impact of electron polarization on scattering off unoriented light nuclei using the unified electroweak theory. Numerical calculations of the scattering cross sections in different electron polarization cases for three 6,7Li and 7Be light nuclei are obtained by directly calculating their multipole form factors, using the formulas provided in Ref. [14]. The results obtained are then compared and/or correlated with the corresponding cross sections for the unpolarized electron scattering case. It has been found that at the scattering angle of $\theta=0^{\circ}$, the longitudinal electron polarization does not explicitly correlate with the weak interaction for all energy scales. Such a correlation is strongly exhibited at nonzero scattering angles ($\theta=5^{\circ}-8^{\circ}$) and high energies above 10 GeV. By comparing the results obtained for stable and unstable nuclei, we found that the nuclear stability significantly influences the contribution of electron polarization to the scattering cross section. This research direction can be extended in many ways, specifically considering various state transitions within a single nucleus or different nuclei. Then, the asymmetry resulting solely from electron polarization and the weak interaction can be figured out. In addition, other interesting results are expected once the nuclear spin orientation and charge-changing weak current are included. Moreover, the multipole expansion technique for the scattering cross section within the unified electroweak theory also suggests a new approach to directly study the EMC effect in a target nucleus using both unpolarized and polarized electrons. More specifically, we calculate the accumulative cross section contributing by separate nucleons inside a certain nucleus and compare it with that made of a similar number of nucleons bound inside the same one. It means that we can investigate the above effect upon only a single nucleus instead of indirectly comparing its scattering cross section with that of a lighter nucleus. These extensive studies will provide further vital information, contributing to a thorough understanding on the nature of the EMC effect in deep inelastic electron-nucleus scattering.
Table 1: Calculated cross-section ratios for 6,7Li and 7Be obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and small scattering angles.

### v2

@@ Abstract
Electron polarization plays a significant role in studies on nuclear scattering. Nevertheless, the development of a comprehensive approach to such a problem remains challenging, particularly at the relativistic electron-energy scale. Herein, we present a theoretical approach to investigate the impact of electron polarization in scattering off unoriented light nuclei, based on the multipole expansion for the scattering cross section at high energies. Numerical calculations for stable 6,7Li and unstable 7Be nuclei show that the longitudinal polarization and weak interaction are not explicitly correlated when electrons scatter at 0 for all energy scales. In contrast, their correlation is strongly exhibited at other scattering angles when energy exceeds 10 GeV. A comparison between stable and unstable nuclei reveals that the stability of nuclei significantly affects the electron polarization contribution to the scattering cross section. This study opens new approaches to the nuclear structure problems, in particular the EMC effect, via deep inelastic electron-nucleus scattering using the unified electroweak theory.
@@
$\sigma_{\text{total}}^{\text{pol.}}(\xi,\xi^{\prime})=\sigma_{\text{total}}^{0}+\sigma_{\text{total}}^{\xi\xi^{\prime}},\vskip-20.0pt$ (1)
where $\xi$ and $\xi^{\prime}$ are the electron polarizations before and after scattering, whose values are equal to +1 or 1-1- 1. The first term in Eq. (1), which does not involve electron polarization, is a sum of
@@
Obviously, the polarized cross sections [Eqs. (3a)–(3c)] depend on the polarization degrees $\xi$ and $\xi^{\prime}$. For simplicity, we assign $\xi\,(\xi^{\prime})=+1$ for the right-handed polarization and $\xi\,(\xi^{\prime})=-1$ for the left-handed one. We then consider all possible variations of electron polarization as following cases. Case 1: both incoming and outgoing electron beams are unpolarized. This case has been studied in detail in our recent study [14]. Case 2: electron polarizations before and after scattering are preserved, i.e., $\xi=\xi^{\prime}$. Case 3: an initially polarized beam becomes unpolarized, i.e., $\xi=+1/-1$ and $\xi^{\prime}=0$ (or $\xi^{\prime}$ is taken over +1 and 1-1- 1). Case 4: an electron beam reverses its initial polarization after scattering, i.e., $\xi=-\xi^{\prime}$.
@@
Fig. 1 shows the calculated $\sigma_{\text{total}}^{\text{pol.}}(1,0)/\sigma_{\text{total}}^{\text{unpol.}}$ ratio for 6Li, whereas Fig. 2 plots the computed $\sigma_{\text{total}}^{\text{pol.}}(1,1)/\sigma_{\text{total}}^{\text{unpol.}}$ one for 7Li. The $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ ratio for 7Be is also displayed in Fig. 3. All the cross section ratios for three considered nuclei are listed in Tab. I. It can be seen that at energies not exceeding 101010 GeV, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ is just half of $\sigma_{\text{total}}^{\text{unpol.}}$, while $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is almost equal to the latter. A slight difference is observed when comparing the cross-section ratios for three nuclei (Tab. I). This can be attributed to the negligible contribution of weak interaction to the polarized terms at such an energy range [13]. When the electron energy goes above 10 MeV, the weak interaction gradually emerges and correlates with electron polarization to make polarized cross sections considerably increase, as seen in Figs. 1–3.
@@
In conclusion, we have performed a systematic study on the impact of electron polarization on scattering off unoriented light nuclei using the unified electroweak theory. Numerical calculations of the scattering cross sections in different electron polarization cases for three 6,7Li and 7Be light nuclei are obtained by directly calculating their multipole form factors, using the formulas provided in Ref. [14]. The results obtained are then compared and/or correlated with the corresponding cross sections for the unpolarized electron scattering case. It has been found that at the scattering angle of $\theta=0^{\circ}$, the longitudinal electron polarization does not explicitly correlate with the weak interaction for all energy scales. Such a correlation is strongly exhibited at nonzero scattering angles ($\theta=5^{\circ}-8^{\circ}$) and high energies above 10 GeV. By comparing the results obtained for stable and unstable nuclei, we found that the nuclear stability significantly influences the contribution of electron polarization to the scattering cross section. This research direction can be extended in many ways, specifically considering various state transitions within a single nucleus or different nuclei. Then, the asymmetry resulting solely from electron polarization and the weak interaction can be figured out. In addition, other interesting results are expected once the nuclear spin orientation and charge-changing weak current are included. Moreover, the multipole expansion technique for the scattering cross section within the unified electroweak theory also suggests a new approach to directly study the EMC effect in a target nucleus using both unpolarized and polarized electrons. More specifically, we calculate the accumulative cross section contributing by separate nucleons inside a certain nucleus and compare it with that made of a similar number of nucleons bound inside the same one. It means that we can investigate the above effect upon only a single nucleus instead of indirectly comparing its scattering cross section with that of a lighter nucleus. These extensive studies will provide further vital information, contributing to a thorough understanding of the nature of the EMC effect in deep inelastic electron-nucleus scattering.
Table 1: Calculated cross section ratios for 6,7Li and 7Be obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and small scattering angles.